# Visual and Memory Dual Adapter for Multi-Modal Object Tracking

### v1

@@ Abstract
Prompt-learning-based multi-modal trackers have achieved promising progress by employing lightweight visual adapters to incorporate auxiliary modality features into frozen foundation models. However, existing approaches often struggle to learn reliable prompts due to limited exploitation of critical cues across frequency and temporal domains. In this paper, we propose a novel visual and memory dual adapter (VMDA) to construct more robust and discriminative representations for multi-modal tracking. Specifically, we develop a simple but effective visual adapter that adaptively transfers discriminative cues from auxiliary modality to dominant modality by jointly modeling the frequency, spatial, and channel-wise features. Additionally, we design the memory adapter inspired by the human memory mechanism, which stores global temporal cues and performs dynamic update and retrieval operations to ensure the consistent propagation of reliable temporal information across video sequences. Extensive experiments demonstrate that our method achieves state-of-the-art performance on the various multi-modal tracking tasks, including RGB-Thermal, RGB-Depth, and RGB-Event tracking. Code and models are available at https://github.com/xuboyue1999/mmtrack.git.
Index Terms:
Multi-modal tracking, prompt learning, visual adapter, frequency domain, memory mechanism.
@@
I Introduction
Visual object tracking (VOT) [1, 2, 3, 4] aims at capturing the target labeled in the first frame in subsequent frames [5] and has widespread applications in autonomous driving [6, 7, 8], embodied artificial intelligence [9, 10], and human-computer interaction [11, 12]. Despite the significant progress achieved by solely RGB-based trackers, they lead to drifting or missing the object in complex scenarios due to the inherent limitations of visible sensors, such as background clutter and low illumination. Therefore, some researchers attempt to enhance the robustness of trackers by introducing auxiliary modalities to develop various multi-modal object tracking tasks, such as RGB-Thermal (RGB-T) [13, 14, 15], RGB-Depth (RGB-D) [16, 17, 18], and RGB-Event (RGB-E) [19, 20].
@@ I Introduction
Multi-modal tracking methods can be broadly divided into two main paradigms. The first is the classic dual-branch architecture [21, 22, 17], typically designed for a specific multi-modal tracking task within a full fine-tuning paradigm. While effective in capturing modality-specific features, these fully fine-tuned approaches often suffer from limitations such as limited dataset availability and high computational cost during training. The second paradigm is based on prompt learning [23, 24, 25], which only fine-tunes lightweight adapters on pre-trained RGB-based trackers, thereby improving both adaptability and training efficiency. Despite the notable progress and impressive performance of prompt-learning-based methods, two critical challenges remain: (1) Frequency cues are essential in multi-modal tasks [26, 27], however, current prompt-learning trackers overlook the significance of the frequency-domain information and fail to effectively bridge the modality gap. We observe that different modalities exhibit distinct and often complementary frequency characteristics due to their sensing principles, as shown in Figure 2. While RGB captures fine-grained high-frequency textures, modalities like Thermal and Depth emphasize low-frequency structural contours. Event data, on the other hand, contains sparse but sharp high-frequency signals representing motion edges. Direct fusion in the spatial domain often leads to sub-optimal performance due to cross-modality noise and misaligned frequency content. While some recent works [28, 29] have begun to explore frequency-aware designs, they mostly rely on manually designed frequency separation schemes, which constrain the model’s ability to fully leverage frequency characteristics in a data-driven manner. (2) Temporal cues are equally crucial in visual object tracking, yet existing trackers [30, 31, 32] lack a full exploration of long-range frame dependencies. These methods often depend on temporal propagation from adjacent frames or confidence-based template updates, which can fail to capture global temporal patterns. As a result, the accumulated temporal errors, especially in scenarios involving occlusion or rapid motion, can significantly degrade tracking robustness.
Figure 2: Illustration of frequency-domain characteristics for RGB-T, RGB-D and RGB-E.
@@
To address these two core challenges, we propose a unified multi-modal tracker based on a dual adapter architecture, comprising a visual adapter and a memory adapter, as illustrated in Figure 1 (b). To facilitate effective modal fusion, we propose a frequency selector that adaptively enhances frequency components for intra-modal feature refinement. In addition, a multi-modal fusion module is designed to generate modality-aware prompts through joint spatial and channel attention, allowing efficient and robust cross-modal interaction.
@@
To model temporal dependencies, we propose the memory adapter inspired by the human memory mechanism [33, 34]. We design a multi-level memory pool to store the temporal cues, consisting of short-term, long-term, and permanent memory, which facilitates the object searching process. We also introduce the memory update and retrieval operations. The update operation refreshes the memory pool with each frame to ensure reliability, while the retrieval operation retrieves the appropriate temporal cues for subsequent tracking. Extensive experiments on mainstream multi-modal tracking benchmarks demonstrate that our tracker outperforms both fully fine-tuned and existing prompt-learning-based methods in terms of accuracy and robustness.
@@
$\bullet$ We propose a unified multi-modal tracker based on the dual adapter architecture, showing superiority on multiple multi-modal tracking tasks, i.e., RGB-T, RGB-D, and RGB-E.
$\bullet$ We develop a lightweight visual adapter that aggregates informative target cues from frequency, spatial, and channel dimensions for multi-modal feature enhancement.
$\bullet$ We present the multi-level memory adapter to effectively store and retrieve the global temporal cues, and adaptively propagate the temporal context among successive frames.
II Related Work
II-A Multi-modal Object Tracking
@@
II-B Temporal Modeling in Visual Tracking
@@
III Method
III-A Preliminary and Notation
@@
$T:\{V\{\textbf{X}_{rgb}^{t},\textbf{X}_{x}^{t}\},V\{\textbf{Z}_{rgb}^{0},\textbf{Z}_{x}^{0},...,\textbf{Z}_{rgb}^{i},\textbf{Z}_{x}^{i}\},\textbf{C}^{t}\}\rightarrow\textbf{B}^{t},$ (2)
@@
Foundation Model. We choose a powerful RGB tracker ODTrack [51] as the foundation model. Given the input search region and template $\textbf{X}_{rgb}$ and $\textbf{Z}_{rgb}$, they first sent into patch embedding layer to obtain 1D tokens $\mathcal{H}_{x}^{0}$ and $\mathcal{H}_{z}^{0}$. These tokens are then concatenated to form the input token $\mathcal{H}^{0}$ = [$\mathcal{H}_{x}^{0}$, $\mathcal{H}_{z}^{0}$]. The input tokens are fed into $L$-layers vision transformer block encoder and the output is passed through a box head to generate the tracking results. The prorogation process can be formulated as follows:
$\mathcal{H}^{l}$ $=\mathcal{E}^{l}(\mathcal{H}^{l-1}),l=1,2,3,...,L,$ (3)
@@ III-C Visual Adapter
III-D Memory Adapter
Memory mechanisms are commonly used in VOT to store temporal tracking cues [51, 30]. To further improve the robustness of multi-modal object tracking, we propose a memory adapter inspired by the human memory system, consisting of short-term memory, long-term memory, and permanent memory. The proposed multi-level memory operates through two key operations: memory update and memory retrieval. In the memory update operation, we use the previous tracking cue token to update all memory levels, while in the memory retrieval operation, the $T$-th temporal tracking cue is retrieved from the memory. Each level holds an $N\!\times\!H$ tensor, where $H\!=\!768$ aligns with the ViT token dimension and $N$ is set to 8, 8, and 3 for short-term, long-term, and permanent memory, respectively. The short-term memory stores tokens from the most recent eight frames, while the long-term and permanent memory banks are hierarchically refreshed following a temporal update protocol.
@@ C $=\textbf{C}_{s}\oplus\textbf{C}_{l}\oplus\textbf{C}_{p},$ (17)
III-E Prediction Head and Loss Function
@@
$\mathcal{L}_{reg}=\sum_{t}\left(\lambda_{1}\mathcal{L}_{1}\left(b_{t},\hat{b_{t}}\right)+\lambda_{2}\mathcal{L}_{GIoU}\left(b_{t},\hat{b_{t}}\right)\right),$ (20)
@@
IV Experiment
IV-A Datasets and Metrics
@@
RGB-T tracking. For RGB-T tracking, we evaluate our method on the latest RGB-T tracking datasets, RGBT234 [14] and LasHeR [13], which are the largest RGB-T tracking dataset with more than 200 testing video sequences of different challenging attributes. We use Precision Rate (PR) and Success Rate (SR) as primary measures and the threshold of center location error is set to 20 pixels. PR represents the ratio of frames $f_{p}$ with center error smaller than a threshold to the total number of frames $N$, which can be calculated as:
@@
IV-B Experimental Settings
When fine-tuning the proposed method, we choose the training sets of LasHeR for RGB-T tracking, DepthTrack for RGB-D tracking, and VisEvent for RGB-E tracking. The proposed method is trained on one NVIDIA RTX 4090 GPU with a batch size of 16. We use the ViT-base trained on MAE as our baseline. The training consists of two stages: in the first stage, we fine-tune the visual adapter and patch-embedding layer for 60 epochs while freezing the other part of the network. In the second stage, we fine-tune the memory adapter on top of stage one for another 60 epochs and freeze the other part except for the patch-embedding layer and visual adapter. Each epoch contains 10,000 samples and we use the AdamW optimizer with a learning rate of $5e^{-4}$.
IV-C Comparison with the State-of-the-Art Methods
We compare the proposed method with two categories of state-of-the-art multi-modal trackers: traditional two-branch methods and prompt-learning methods. The former includes methods specifically designed for a particular type of multi-modal tracking that fully fine-tune the two-branch network for both modalities. The latter encompasses methods designed for general multi-modal tracking that only fine-tune the adapters and patch-embedding layers.
RGB-T tracking. The comparison results for RGB-T tracking are shown in Table I, where the proposed method outperforms all existing methods in terms of precision rate and success rate. Specifically, it exceeds the second-best method, TaTrack [55], on the RGBT234 dataset by 4.7% and 4.5% gains in PR and SR, respectively. On the LasHeR dataset, the proposed method achieves improvements of 2.4% and 0.8% in PR and SR compared to BaT [56]. Additionally, compared to the latest two-branch tracker, STMT [57], the proposed method shows enhancements of 5.2% and 3.4% in PR and SR, respectively.
TABLE I: Comparison between the proposed method and the state-of-the-art trackers on RGB-T datasets. The best results are highlighted in bold. The performance is evaluated in terms of Precision Rate (PR) and Success Rate (SR).
Methods Years RGBT234 LasHeR Param FPS
@@ IV-C Comparison with the State-of-the-Art Methods
TABLE II: Comparison between the proposed method and the state-of-the-art trackers on RGB-D datasets. The best results are highlighted in bold. The performance is evaluated in terms of precision (Pre), recall (Re), F-score(F) on DepthTrack and EAO, accuracy(A), robustness(R) on VOT-RGBD22.
Methods Years DepthTrack VOT-RGBD22
@@
TABLE III: Comparison between the proposed method and the state-of-the-art trackers on RGB-E datasets. The best results are highlighted in bold. The performance is evaluated in terms of precision rate (PR), success rate (SR).
Methods Years VisEvent
@@
Figure 6: Precision scores of different attributes on the VisEvent test set.
RGB-D tracking. The experimental results for RGB-D tracking are presented in Table II. The proposed method achieves 63.6%, 66.3%, and 64.9% in precision, recall, and F-score on the DepthTrack dataset, and 77.3%, 82.1%, and 93.3% in EAO, accuracy, and robustness on the VOT-RGBD22 dataset, respectively. These results demonstrate that the proposed tracker significantly outperforms existing methods across all metrics. Notably, it surpasses the second-best method, SDSTrack [23], with gains of 5% in robustness and 4.5% in EAO.
RGB-E tracking. We compare the proposed method with the state-of-the-art RGB-E trackers. As shown in Table III, the proposed method outperforms the best prompt-learning method, OneTracker [25] with improvements of 3.6% in PR and 1.8% in SR. Additionally, the proposed method achieves gains of 9.3% in PR and 6.1% in SR over SwinEFT [20], which is specifically designed for RGB-E tracking. These results demonstrate that our tracker also exhibits strong generalization capability in various tracking tasks.
TABLE IV: Component analysis on multi-modal tracking datasets. Visual represents the proposed visual adapter, and memory represents the memory adapter.
@@
TABLE V: Comparison of visual adapters on multi-modal tracking datasets.
@@
TABLE VI: Comparison of tunable parameters on the LasHeR dataset, the performance is evaluated in terms of precision rate (PR) and success rate (SR).
@@
IV-D Ablation Studies.
1) Effectiveness of Components: We conduct an ablation study to evaluate key components of our method. As shown in Table IV, the baseline (Row 1) fuses RGB and auxiliary modalities via element-wise addition with fine-tuned patch embedding, yielding initial results. Replacing element-wise addition with our visual adapter (Row 2) significantly improves performance, notably on LasHeR (+5.9% PR and +4.7% SR). On DepthTrack, this change lifts the F-score from 0.598 to 0.626 (+2.8 points), indicating that frequency-aware recalibration benefits sequences where depth maps are noisy but still contain high-frequency edge cues. Integrating the multi-level memory into the baseline (Row 3) enhances robustness across all datasets. The gain is most evident on VisEvent, where SR rises from 0.605 to 0.618, confirming that temporal aggregation helps smooth the bursty event stream and recover from momentary information loss. The full method (Row 4), combining both components, achieves the highest performance gains, demonstrating their synergistic effectiveness in boosting tracking accuracy.
@@
TABLE VII: Effectiveness of visual adapter on the LasHeR dataset, the performance is evaluated in terms of precision rate (PR) and success rate (SR).
@@
To further validate the effectiveness of the proposed visual adapter, we conduct a comprehensive ablation study to investigate the contributions of its internal modules. The experimental results on the LasHeR dataset are presented in Table VII. We decompose the Visual Adapter into three components: spatial, channel, and frequency modules. The first row displays baseline performance without the Visual Adapter, achieving PR (Precision Rate) and SR (Success Rate) of 0.659 and 0.518, respectively. When incorporating spatial information processing (second row), both metrics increased by 3.1% and 2.6%. The subsequent integration of channel information (third row) yielded additional improvements of 0.8% and 0.5%. The complete Visual Adapter (final row), incorporating all three modules, further enhanced performance by 2% and 1.6% in PR and SR, respectively. The experiments conclusively validate the effectiveness of each component in the Visual Adapter and demonstrate the necessity of integrating frequency information.
@@
TABLE VIII: Effectiveness of memory adapter on the LasHeR dataset, the performance is evaluated in terms of precision rate (PR) and success rate (SR).
@@
To further investigate the effectiveness of the Memory Adapter, the experimental results on LasHeR is shown in Table VIII. The baseline method, employing ODTrack’s adjacent-frame propagation approach, achieves PR and SR scores of 0.659 and 0.518, respectively. Incorporating short-term memory improves these metrics by 1.3% and 0.9%, while subsequent integration of long-term memory yields additional gains of 0.8% and 1.0%. The complete Memory Adapter ultimately attains PR and SR scores of 0.689 and 0.545. These results demonstrate that our memory adapter more effectively utilizes temporal information compared to adjacent-frame propagation methods, while also confirming the rationality of our design.
IV-E Parameter Analysis.
TABLE IX: Memory configuration study on LasHeR. “S/L/P” denotes the number of tokens in Short-, Long- and Permanent-term banks, respectively.
@@
To investigate how memory size and read-out strategy affect tracking performance, we conduct a parameter study on the LasHeR dataset, as shown in Table IX. Starting from the strong visual-adapter baseline that has no memory at all (PR = 71.8%, SR = 56.5%), we first enable the memory module while keeping a large long-term bank of 8/8/8 tokens and our similarity-based retrieval. This configuration boosts performance to 72.2% and 56.8%, showing that simply giving the model more historical snapshots already helps it recover appearance information lost during occlusion. Shrinking the bank to 3/3/3 tokens, while still using similarity retrieval, reduces the gain to 72.0% and 56.7%, which implies that an undersized memory cannot fully capture the appearance diversity of the long LasHeR sequences.
@@
IV-F Attribute Analysis.
@@
We perform an analysis of various challenging attributes on VisEvent including Illumination Variation, Over Exposure, Background Object Motion, Rotation, Camera Motion, Deformation, Full occlusion, Low illumination, Out of view, Partial Occlusion, Viewpoint Change, Scale Variation, Background Clutter, Motion Blur, Aspect Ratio-Change, Fast Motion, No Motion. As in Figure 6. The Figure clearly shows that our tracker encloses a consistently larger area than all competing methods, indicating broad improvements rather than isolated wins on a few attributes. The gain is most pronounced on full occlusion and overexposure, where the RGB stream is unreliable but the event modality still captures edge activity; the visual adapter dynamically up-weights the latter, sustaining stable localization after re-appearance. We also observe non-trivial margins under fast motion and low illumination. This suggests that the frequency-aware recalibration mitigates motion blur and intensity saturation by selectively emphasizing high-frequency cues from events while suppressing noisy low-frequency components in RGB. This improvement is attributed to the visual adapter, which selectively leverages multi-modal information, thereby enhancing robustness and adaptability in complex environments.
IV-G Comparison of Tunable Parameter.

### v2

# Learning Frequency and Memory-Aware Prompts for Multi-Modal Object Tracking

@@ Abstract
Prompt-learning–based multi-modal trackers have made strong progress by using lightweight visual adapters to inject auxiliary-modality cues into frozen foundation models. However, they still underutilize two essentials: modality-specific frequency structure and long-range temporal dependencies. We present Learning Frequency and Memory-Aware Prompts, a dual-adapter framework that injects lightweight prompts into a frozen RGB tracker. A frequency-guided visual adapter adaptively transfers complementary cues across modalities by jointly calibrating spatial, channel, and frequency components, narrowing the modality gap without full fine-tuning. A multilevel memory adapter with short, long, and permanent memory stores, updates, and retrieves reliable temporal context, enabling consistent propagation across frames and robust recovery from occlusion, motion blur, and illumination changes. This unified design preserves the efficiency of prompt learning while strengthening cross-modal interaction and temporal coherence. Extensive experiments on RGB-Thermal, RGB-Depth, and RGB-Event benchmarks show consistent state-of-the-art results over fully fine-tuned and adapter-based baselines, together with favorable parameter efficiency and runtime. Code and models are available at https://github.com/xuboyue1999/mmtrack.git.
keywords:
Multi-modal tracking , prompt learning , frequency-guided fusion , memory mechanism , temporal modeling.
journal: Pattern Recognition
[…]
1 Introduction
Visual object tracking (VOT) [1, 2, 3, 4] aims to localize the target annotated in the first frame throughout subsequent frames [5] and underpins a wide range of applications in autonomous driving [6, 7, 8], embodied artificial intelligence [9, 10], and human–computer interaction [11, 12]. Despite the significant progress of RGB-only trackers, they can still drift or miss the target in complex scenarios due to the inherent limitations of visible-spectrum sensors, such as background clutter and low illumination. To enhance robustness, recent works introduce auxiliary modalities and develop multi-modal tracking tasks, including RGB-Thermal (RGB-T) [13, 14, 15], RGB-Depth (RGB-D) [16, 17, 18], and RGB-Event (RGB-E) [19, 20].
@@ 1 Introduction
Multi-modal tracking methods generally fall into two paradigms. The first is the classic dual-branch architecture [21, 22, 17], typically tailored to a specific multi-modal setting and trained end-to-end with full fine-tuning. While effective at capturing modality-specific representations, such methods are constrained by limited training data and high computational cost. The second paradigm adopts prompt learning [23, 24, 25], which fine-tunes lightweight adapters on frozen RGB-based backbones, improving adaptability and training efficiency. Despite notable progress, two critical gaps remain. (1) Frequency cues are essential in multi-modal tracking [26, 27], yet current prompt-learning trackers underuse frequency-domain information and thus struggle to bridge the modality gap. Different modalities exhibit complementary frequency characteristics due to their sensing principles (Fig. 2): RGB encodes fine-grained high-frequency textures; Thermal and Depth emphasize low-frequency structural contours; Event streams carry sparse, sharp high-frequency motion edges. Direct spatial-domain fusion is therefore suboptimal owing to cross-modality noise and misaligned frequency content. Although recent efforts explore frequency-aware designs [28, 29], they largely rely on hand-crafted frequency separation, limiting data-driven exploitation. (2) Temporal cues are equally crucial, yet existing trackers [30, 31, 32] typically restrict modeling to adjacent-frame propagation or confidence-based template updates, which fail to capture long-range dependencies; errors accumulate under occlusion, fast motion, or appearance changes, degrading robustness.
Figure 2: Illustration of frequency-domain characteristics for RGB-T, RGB-D and RGB-E. The second to fourth rows show the magnitude map, the low-frequency visualization, and the high-frequency visualization, respectively. Red boxes indicate the most informative regions in the low-frequency domain, while green boxes highlight the most informative regions in the high-frequency domain.
@@
To address these two core challenges, we propose a unified multi-modal tracker that learns frequency and memory-aware prompts via a dual-adapter architecture comprising a frequency-guided visual adapter and a multi-level memory adapter (Fig. 1(b)). For effective cross-modal fusion, the visual adapter introduces a frequency selector that adaptively emphasizes informative subbands to refine intra-modal features, and a multi-modal fusion module that produces modality-aware prompts through joint spatial and channel attention, thereby narrowing the modality gap without full fine-tuning.
@@
To model temporal dependencies, the memory adapter is inspired by the human memory mechanism [33, 34]. It maintains a multi-level memory pool (short-term, long-term, and permanent) to store global temporal cues, together with update and retrieval operations that refresh the memory per frame and select reliable cues for subsequent tracking. This design learns memory-aware prompts that propagate consistent temporal context across frames and recover from occlusion, fast motion, and illumination changes. Extensive experiments on mainstream multi-modal tracking benchmarks demonstrate that our tracker outperforms fully fine-tuned and prompt-learning baselines in both accuracy and robustness, while preserving the efficiency of adapter-based tuning.
@@
$\bullet$ We present a unified dual-adapter framework that learns frequency and memory-aware prompts for multi-modal tracking, showing consistent gains on RGB-T, RGB-D, and RGB-E tasks.
$\bullet$ We develop a lightweight frequency-guided visual adapter that aggregates informative cues across frequency, spatial, and channel dimensions to produce modality-aware prompts and enhance cross-modal fusion.
$\bullet$ We propose a multi-level memory adapter that stores and retrieves global temporal cues with update and retrieval operations, enabling adaptive propagation of temporal context along video sequences.
2 Related Work
2.1 Multi-modal Object Tracking
@@
2.2 Temporal Modeling in Visual Tracking
@@
3 Method
3.1 Preliminary and Notation
@@
$T:\{V\{\textbf{X}_{rgb}^{t},\textbf{X}_{x}^{t}\},V\{\textbf{Z}_{rgb}^{0},\textbf{Z}_{x}^{0},.,\textbf{Z}_{rgb}^{i},\textbf{Z}_{x}^{i}\},\textbf{C}^{t}\}\rightarrow\textbf{B}^{t},$ (2)
@@
Foundation Model. We choose a powerful RGB tracker ODTrack [51] as the foundation model. Given the input search region and template $\textbf{X}_{rgb}$ and $\textbf{Z}_{rgb}$, they first sent into patch embedding layer to obtain 1D tokens $\mathcal{H}_{x}^{0}$ and $\mathcal{H}_{z}^{0}$. These tokens are then concatenated to form the input token $\mathcal{H}^{0}$ = $\mathcal{H}_{x}^{0}$, $\mathcal{H}_{z}^{0}$]. The input tokens are fed into $L$-layers vision transformer block encoder and the output is passed through a box head to generate the tracking results. The prorogation process can be formulated as follows:
$\mathcal{H}^{l}$ $=\mathcal{E}^{l}(\mathcal{H}^{l-1}),l=1,2,3,.,L,$ (3)
@@ 3.3 Visual Adapter
3.4 Memory Adapter
Memory mechanisms are commonly used in VOT to store temporal tracking cues [51, 30]. To further improve the robustness of multi-modal object tracking, we propose a memory adapter inspired by the human memory system, consisting of short-term memory, long-term memory, and permanent memory. The proposed multi-level memory operates through two key operations: memory update and memory retrieval. In the memory update operation, we use the previous tracking cue token to update all memory levels, while in the memory retrieval operation, the $T$-th temporal tracking cue is retrieved from the memory. Each level holds an $N\!\times\!H$ tensor, where H=768H\!=\!768 aligns with the ViT token dimension and $N$ is set to 8, 8, and 3 for short-term, long-term, and permanent memory, respectively. The short-term memory stores tokens from the most recent eight frames, while the long-term and permanent memory banks are hierarchically refreshed following a temporal update protocol.
@@ C $=\textbf{C}_{s}\oplus\textbf{C}_{l}\oplus\textbf{C}_{p},$ (17)
3.5 Prediction Head and Loss Function
@@
[…]
@@
4 Experiment
4.1 Datasets and Metrics
@@
RGB-T tracking. For RGB-T tracking, we evaluate our method on the latest RGB-T tracking datasets, RGBT234 [14] and LasHeR [13], which are the largest RGB-T tracking datasets with more than 200 testing video sequences of different challenging attributes. We use Precision Rate (PR) and Success Rate (SR) as primary measures and the threshold of center location error is set to 20 pixels. PR represents the ratio of frames $f_{p}$ with center error smaller than a threshold to the total number of frames $N$, which can be calculated as:
@@
4.2 Experimental Settings
When fine-tuning the proposed method, we choose the training sets of LasHeR for RGB-T tracking, DepthTrack for RGB-D tracking, and VisEvent for RGB-E tracking. The proposed method is trained on one NVIDIA RTX 4090 GPU with a batch size of 16. We use the ViT-base trained on MAE as our baseline. The training consists of two stages: in the first stage, we fine-tune the visual adapter and patch-embedding layer for 60 epochs while freezing the other part of the network. In the second stage, we fine-tune the memory adapter on top of stage one for another 60 epochs and freeze the other part except for the patch-embedding layer and visual adapter. Each epoch contains 10,000 samples and we use the AdamW optimizer with a learning rate of $5e^{-4}$. In addition, our method contains a total of 98.9M parameters, introducing 7.3M more than the baseline, with an additional computational cost of 1 GFLOPS.
4.3 Comparison with the State-of-the-Art Methods
We compare the proposed method with two categories of state-of-the-art multi-modal trackers: traditional two-branch methods and prompt-learning methods. The former includes methods specifically designed for a particular
Table 1: Comparison between the proposed method and the state-of-the-art trackers on RGB-T datasets. The best results are highlighted in bold. The performance is evaluated in terms of Precision Rate (PR) and Success Rate (SR).
Methods Publication RGBT234 LasHeR Param FPS
@@ 4.3 Comparison with the State-of-the-Art Methods
type of multi-modal tracking that fully fine-tunes the two-branch network for both modalities. The latter encompasses methods designed for general multi-modal tracking that only fine-tune the adapters and patch-embedding layers.
RGB-T tracking. The comparison results for RGB-T tracking are shown in Table 1, where the proposed method outperforms all existing methods in terms of precision rate and success rate. Specifically, it exceeds the second-best method, TaTrack [59], on the RGBT234 dataset by 4.7% and 4.5% gains in PR and SR, respectively. On the LasHeR dataset, the proposed method achieves improvements of 2.4% and 0.8% in PR and SR compared to BaT [60]. Additionally, compared to the latest two-branch tracker, STMT [56], the proposed method shows enhancements of 5.2% and 3.4% in PR and SR, respectively.
Table 2: Comparison between the proposed method and the state-of-the-art trackers on RGB-D datasets. The best results are highlighted in bold. The performance is evaluated in terms of precision (Pre), recall (Re), F-score(F) on DepthTrack and EAO, accuracy(A), robustness(R) on VOT-RGBD22.
Methods Publication DepthTrack VOT-RGBD22
@@
Table 3: Comparison between the proposed method and the state-of-the-art trackers on RGB-E datasets. The best results are highlighted in bold. The performance is evaluated in terms of precision rate (PR) and success rate (SR).
Methods Publication VisEvent
@@
RGB-D tracking. The experimental results for RGB-D tracking are presented in Table 2. The proposed method achieves 63.6%, 66.3%, and 64.9% in precision, recall, and F-score on the DepthTrack dataset, and 77.3%, 82.1%, and 93.3% in EAO, accuracy, and robustness on the VOT-RGBD22 dataset, respectively. These results demonstrate that the proposed tracker significantly outperforms existing methods across all metrics. Notably, it surpasses the second-best method, SDSTrack [23], with gains of 5% in robustness and 4.5% in EAO.
RGB-E tracking. We compare the proposed method with the state-of-the-art RGB-E trackers. As shown in Table 3, the proposed method outperforms the best prompt-learning method, OneTracker [25] with improvements of 3.6% in PR and 1.8% in SR. Additionally, the proposed method achieves gains of 9.3% in PR and 6.1% in SR over SwinEFT [20], which is specifically designed for RGB-E tracking. These results demonstrate that our tracker also exhibits strong generalization capability in various tracking tasks.
Table 4: Component analysis on multi-modal tracking datasets. Visual represents the proposed visual adapter, and memory represents the memory adapter.
@@
4.4 Ablation Studies.
1) Effectiveness of Components: We conduct an ablation study to evaluate key components of our method. As shown in Table 4, the baseline (Row 1) fuses RGB and auxiliary modalities via element-wise addition with fine-tuned patch embedding, yielding initial results. Replacing element-wise addition with our visual adapter (Row 2) significantly improves performance, notably on LasHeR (+5.9% PR and +4.7% SR). On DepthTrack, this change lifts the F-score from 0.598 to 0.626 (+2.8 points), indicating that frequency-aware recalibration benefits sequences where depth maps are noisy but still contain high-frequency edge cues. Integrating the multi-level memory into the baseline (Row 3) enhances robustness across all datasets. The gain is most evident on VisEvent, where SR rises from 0.605 to 0.618, confirming that temporal aggregation helps smooth the bursty event stream and recover from momentary information loss. The full method (Row 4), combining both components, achieves the highest performance gains, demonstrating their synergistic effectiveness in boosting tracking accuracy.
Table 5: Comparison of visual adapters on multi-modal tracking datasets.
@@
Table 6: Comparison of tunable parameters on the LasHeR dataset, the performance is evaluated in terms of precision rate (PR) and success rate (SR).
@@
Table 7: Effectiveness of visual adapter on the LasHeR dataset, the performance is evaluated in terms of precision rate (PR) and success rate (SR).
@@
To further validate the effectiveness of the proposed visual adapter, we conduct a comprehensive ablation study to investigate the contributions of its internal modules. The experimental results on the LasHeR dataset are presented in Table 7. We decompose the Visual Adapter into three components: spatial, channel, and frequency modules. The first row displays baseline performance without the Visual Adapter, achieving PR (Precision Rate) and SR (Success Rate) of 0.659 and 0.518, respectively. When incorporating spatial information processing (second row), both metrics increased by 3.1% and 2.6%. The subsequent integration of channel information (third row) yielded additional improvements of 0.8% and 0.5%. The complete Visual Adapter (final row), incorporating all three modules, further enhanced performance by 2% and 1.6% in PR and SR, respectively. The experiments conclusively validate the effectiveness of each component in the Visual Adapter and demonstrate the necessity of integrating frequency information.
@@
Table 8: Effectiveness of memory adapter on the LasHeR dataset, the performance is evaluated in terms of precision rate (PR) and success rate (SR).
@@
To further investigate the effectiveness of the Memory Adapter, the experimental results on LasHeR are shown in Table 8. The baseline method, employing ODTrack’s adjacent-frame propagation approach, achieves PR and SR scores of 0.659 and 0.518, respectively. Incorporating short-term memory improves these metrics by 1.3% and 0.9%, while subsequent integration of long-term memory yields additional gains of 0.8% and 1.0%. The complete Memory Adapter ultimately attains PR and SR scores of 0.689 and 0.545. These results demonstrate that our memory adapter more effectively utilizes temporal information compared to adjacent-frame propagation methods, while also confirming the rationality of our design.
4.5 Parameter Analysis.
Table 9: Memory configuration study on LasHeR. “S/L/P” denotes the number of tokens in Short-, Long- and Permanent-term banks, respectively.
@@
To investigate how memory size and read-out strategy affect tracking performance, we conduct a parameter study on the LasHeR dataset, as shown in Table 9. Starting from the strong visual-adapter baseline that has no memory at all (PR = 71.8%, SR = 56.5%), we first enable the memory module while keeping a large long-term bank of 8/8/8 tokens and our similarity-based retrieval. This configuration boosts performance to 72.2% and 56.8%, showing that simply giving the model more historical snapshots already helps it recover appearance information lost during occlusion. Shrinking the bank to 3/3/3 tokens, while still using similarity retrieval, reduces the gain to 72.0% and 56.7%, which implies that an undersized memory cannot fully capture the appearance diversity of the long LasHeR sequences.
@@
Figure 6: Precision scores of different attributes on the VisEvent.
4.6 Attribute Analysis.
@@
We perform an analysis of various challenging attributes on VisEvent including Illumination Variation, Over Exposure, Background Object Motion, Rotation, Camera Motion, Deformation, Full Occlusion, Low Illumination, Out of View, Partial Occlusion, Viewpoint Change, Scale Variation, Background Clutter, Motion Blur, Aspect Ratio-Change, Fast Motion, and No Motion. As in Figure 6. The Figure clearly shows that our tracker encloses a consistently larger area than all competing methods, indicating broad improvements rather than isolated wins on a few attributes. The gain is most pronounced on full occlusion and overexposure, where the RGB stream is unreliable but the event modality still captures edge activity; the visual adapter dynamically up-weights the latter, sustaining stable localization after re-appearance. We also observe non-trivial margins under fast motion and low illumination. This suggests that the frequency-aware recalibration mitigates motion blur and intensity saturation by selectively emphasizing high-frequency cues from events while suppressing noisy low-frequency components in RGB. This improvement is attributed to the visual adapter, which selectively leverages multi-modal information, thereby enhancing robustness and adaptability in complex environments.
4.7 Comparison of Tunable Parameter.